Errata to Mathematics for the Analysis of Algorithms, third edition (last updated 26 November 2024)

Page 1, eq. (1.1) integer n integer n ≥ 0 Page 16, in (2.27)

∑_{k=1}^{n−3} ∑_{k=0}^{n−3}

Page 25, after the theorem [Further results on such recurrences have been obtained by T. R. Walsh, Information Processing Letters 19 (1984), 203–208; S. Kapoor and E. M. Reingold, J. Math. Analysis and Applications 109 (1985), 591–604.] Page 25, in (2.64) A_{h+1} A_{h+1}(z) Page 26, line 6 we were lead we were led Page 48, line 2 from the bottom in equation (4.22) in (4.22) Page 77–80, throughout the bibliography adding several commas would make the format more consistent Page 78, reference [Erdős 59] 385–387 290–297 Page 79, line 25 leurs relation leurs relations Page 80, bottom line 5(1) 5(3) Page 87, line 7 (1 1 Page 90, line 8 from the bottom 0676 0739 Page 94, add a new (one-line) paragraph Curiously, (D19) is wrong! Page 104, line 14 from the bottom z_{1} … z_{n−k} z_{2} … z_{n−k} Page 106, bottom of page [Another solution to this problem, which leads to a generating function, has been published by P. V. Poblete, Journal of Algorithms 4 (1983), 388–393.] Page 132, line 26 of the right-hand column Louis Xavier Joseph Jean Gustave Nicolas